Wood paneling can be ordered in thickness of 1/16,1/8,1/4, or3/8inch. The random variable


X is the total thickness of paneling in two orders. List the elements in the sample


space.

1
Expert's answer
2022-03-02T17:05:49-0500

Since all wood paneling has different thickness, then there are "{4 \choose 2}+4=10" different combinations of 2 wood paneling. Let's find total thickness for every of them

"{\frac 1 {16}}+{\frac 1 {16}}={\frac 1 8}"

"{\frac 1 {16}}+{\frac 1 {8}}={\frac 3 {16}}"

"{\frac 1 {16}}+{\frac 1 4}={\frac 5 {16}}"

"{\frac 1 {16}}+{\frac 3 8}={\frac 7 {16}}"

"{\frac 1 8}+{\frac 1 8}={\frac 1 4}"

"{\frac 1 8}+{\frac 1 4}={\frac 3 8}"

"{\frac 1 8}+{\frac 3 8}={\frac 1 2}"

"{\frac 1 4}+{\frac 1 4}={\frac 1 2}"

"{\frac 1 4}+{\frac 3 8}={\frac 5 8}"

"{\frac 3 8}+{\frac 3 8}={\frac 3 4}"

So, X = {"{\frac 1 8},{\frac 3 {16}},{\frac 5 {16}},{\frac 7 {16}},{\frac 1 4},{\frac 3 8},{\frac 1 2},{\frac 5 8},{\frac 3 4}"}. X measures in inches